You bought a stock at $67 and want to use a protective put strategy that consists in buying a put to protect against losses on the stock position. Puts with a strike K=62 are quoted with a bid of $1.08 and an ask of $1.21. What is the maximum loss on the entire strategy?

{Enter the positive number of the loss (do not use a negative sign in your answer) with 2 decimals. For example, if you compute the profit/loss to be -34.56. Then, the loss is 34.56. Enter your answer as 34.56}
Answer #1

You bought a stock at $67 and want to use a protective put strategy that consists in buying a put to protect against losses on the stock position. Puts with a strike K=62 are quoted with a bid of $1.08 and an ask of $1.21. What is the maximum loss on the entire strategy?

Purchase price of the stock = $67

Strike price = $62

We buy put option at the ask price of $1.21

The maximum loss in a protective put occurs when the stock expires at the strike price.

When St = $62, the loss from long stock = 67 - 62 = $5

When St = $62, the put option expires worthless and we would lose the entire premium of $1.21

The maximum loss = 5 + 1.21

The maximum loss = $6.21

Can you please upvote? Thank You :-)